BA Corp is issuing a 10-year bond with a coupon rate of 6.00 percent. The interest rate for similar bonds is currently 4.00 percent. Assuming annual payments, what is the value of the bond? (Round answer to 2 decimal places, e.g. 15.25.)

Value of bond $....?
Answer #1

Annual coupon=1000*6%=60

Hence value of bond=Annual coupon*Present value of annuity factor(4%,10)+$1000*Present value of discounting factor(4%,10)

=60*8.110895779+$1000*0.675564168

=$1162.22(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=60[1-(1.04)^-10]/0.04

=60*8.110895779

2.Present value of discounting factor=1000/1.04^10

=1000*0.675564168